OpenStudy (anonymous):

Create your own polynomial with a degree greater than 2. Find the zeros of the function. could someone walk me through on how to do this? I don't get what "find the zeros" means..

OpenStudy (perl):

mathematicians use different words for the same thing zeros = roots = x intercepts

OpenStudy (perl):

I tend to use the word roots more than 'zeros' so we can look at a factored form a polynomial, in terms of its 'roots' y = a ( x - r1) (x-r2) (x-r3)... (x-rn) , where r1,r2,..rn are roots

OpenStudy (perl):

so you want a polynomial of degree greater than 2, so you will need at least 3 factors y = a ( x - r1 ) ( x - r2) ( x - r3) , now pick any values for r1, r2, r3 and you can let 'a' equal 1 , since thats easiest

OpenStudy (anonymous):

sorry, i still don't really get what you're saying :(

OpenStudy (anonymous):

i had another question that said "Using complete sentences, explain how to find the zeros of the function f(x) = 2^3 – 9x + 3." i wrote that you would try to find the x values in this to check if they came out to 0, is that right?

OpenStudy (anonymous):

function f(x) = 2^3 - 9x + 3.

OpenStudy (perl):

is there an x missing

OpenStudy (anonymous):

f(x) = 2x^3 - 9x + 3.

OpenStudy (perl):

ok, thats a cubic polynomial

OpenStudy (perl):

it is difficult to factor these generally, but sometimes it has nice x intercepts (roots) which I find by graphing it

OpenStudy (perl):

alternatively you can use rational roots theorem, but this theorem won't guarantee you rational roots exist

OpenStudy (perl):

the theorem says *if* rational roots exist, then the roots will have such and such form

OpenStudy (perl):

if f(p/q) = 0 , then p divides the constant coefficient and q divides the coefficient of the degree term

OpenStudy (perl):

and check out wolfram, it does a lot of the tedious stuff. here is what i will submit into wolfram "solve 2^3 - 9x + 3= 0 " (this gives us the roots)

OpenStudy (perl):

Wolfram gives me one real rational root x = 11/9 http://www.wolframalpha.com/input/?i=solve+2^3+-+9x+%2B+3%3D+0

OpenStudy (perl):

so now you can do synthetic/long division and factor out (x -11/9) ,

OpenStudy (perl):

alternatively, you could use something called the cubic formula, but its probably outside the scope of your class or homework

OpenStudy (perl):

and it would be overkill in this example to use cubic formula